OpenStudy (anonymous):

1. Given f(x) = x + 5 and g(x) = 11x - 5, what is f(x) + g(x)? 2. Given f(x) = x + 9 and g(x) = 5x - 5, what is f(3) + g(- 3)? 3. Given f(x) = - 8x + 5 and g(x) = 10x + 2, what is g(x) - f(x)? 4. Given f(x) = 3x - 3 and g(x) = 2x , what is f(4) - g(2)? 5. Given f(x) = 6x + 2 and g(x) = x^3 , what is f(x) times g(x)? 6. Given f(x) = 4x^7 and g(x) = 6x^2 , what is f(x) divided by g(x)?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (anonymous):

@4everaddicted2anime

OpenStudy (anonymous):

1. = 12x 2. = -27 3. = 2x - 3 4. = 5 5. = 6x^4 + 2x^3 6. = 2x^5/3

OpenStudy (4everaddicted2anime):

Number one is correct

OpenStudy (4everaddicted2anime):

For number two you have \[3+9+5(-3)-5\] Multiply -3 and 5 to get \[3+9-15-5\] Combine lilke terms to get \[12-20=-8\]

OpenStudy (anonymous):

the ansers for 2. are ether 27 0r -27 3 0r -3

OpenStudy (4everaddicted2anime):

For number three you have\[10x+2--8x+5\] Combine like terms to get\[18x+7\]

OpenStudy (4everaddicted2anime):

Then I don't know about number two

OpenStudy (anonymous):

ok thats fine

OpenStudy (4everaddicted2anime):

Number 4 is correct

OpenStudy (4everaddicted2anime):

Number 5 is correct

OpenStudy (4everaddicted2anime):

And Number 6 is correct

OpenStudy (4everaddicted2anime):

7. find f(g(10)) if f(x) = 2x and g(x) = 4x - 25 8. find f(g(x)) if f(x) = 5x + 9 and g(x) = 3 - 4x here are a few more to check 7. = 45 8. = 24 - 20x Here is what we have for number seven: \[2[4(10)-25]=?\] Multiply 4 and 10 to get\[2[40-25]=?\] Simplify everything inside the brakets to get\[2(15)=?\] 2 times 15 is 30. Therefore the answer to number 7 is 30.

OpenStudy (4everaddicted2anime):

Number 8 is correct

OpenStudy (anonymous):

thanks

OpenStudy (4everaddicted2anime):

yw